OpenStudy (anonymous):

how do you factor 4asquared-4ab+bsquared-csquared?

OpenStudy (anonymous):

\[4a^2-4ab+b^2-c^2\] like that ?

OpenStudy (inkyvoyd):

Ok, let me explain agin.

OpenStudy (anonymous):

yup

OpenStudy (inkyvoyd):

\((ax-b)^2=a^2x^2-abx+b^2\) right?

OpenStudy (anonymous):

first three terms are a perfect square ,namely \[4a^2-4ab+b^2=(2a-b)^2\]

OpenStudy (anonymous):

what do you mean?

OpenStudy (inkyvoyd):

Hannah, are you talking to me, or @satellite73 ?

OpenStudy (anonymous):

satellite73

OpenStudy (inkyvoyd):

\((ax−b)^2=a^2x^2−abx+b^2\) look at \(4a^2−4ab+b^2\) it equals \((2^2)a^2−(2*2)ab+b^2\)

OpenStudy (anonymous):

inkyvoyd i am sure can explain better than i can. but what i meant by "perfect square" is something that can be written as \((\text{expression})^2\)

OpenStudy (anonymous):

@satellite73 i see how you got the 2a, but i dont understand where you got the b from, did you get it from the -4ab? or the +b^2?

OpenStudy (inkyvoyd):

hannah, you know that \((a+b)^2=a^2+2ab+b^2\) right?

OpenStudy (anonymous):

@inkyvoyd yes

OpenStudy (inkyvoyd):

Ok, then what does \((2a+b)^2\) equal?

OpenStudy (anonymous):

4a+6ab+b^2

OpenStudy (anonymous):

I used the FOIL method

OpenStudy (inkyvoyd):

Careful.

OpenStudy (anonymous):

Whoooops. i meant 4ab not 6ab!!!!!!

OpenStudy (inkyvoyd):

4a^2 though, too.

OpenStudy (anonymous):

o yea theres 2 a's

OpenStudy (inkyvoyd):

Let me try to offer another explanation. \((x+y)^2=x^2+2xy+y^2\) in this case, \(4a^2+4ab+b^2\), 2a=y, and b=y

OpenStudy (anonymous):

wait but i have a question, wh dod you grab the 1st 3 terms?? thats what i dont understand, the square root theory thing is x^2-y^2=(x+y)(x-y), and theres only 2 terms, x and y. but u grabbed the 1st 3 terms and factored

OpenStudy (inkyvoyd):

remember, x^2 is (2a)^2, or 2*2*a*a,

OpenStudy (inkyvoyd):

Alright, look at this again - let me rewrite it \(4a^2+4ab+b^2\) \((2a)^2+2*(2a)b+b^2\) 2a->x b->y

OpenStudy (inkyvoyd):

\((x)^2+2(x)y+y^2\)

OpenStudy (anonymous):

OH! so it's a trinomial!!!!!!!!!!!!!!!! the x^2 is the 4a^2 the 2xy is the 4ab and the b^2 is the y^2!

OpenStudy (inkyvoyd):

Yesh.

OpenStudy (anonymous):

ok, so the factorization of that trinomial would be?

OpenStudy (inkyvoyd):

\((x)^2+2(x)y+y^2=x^2+2xy+y^2=(x+y)^2\)

OpenStudy (anonymous):

yes but in this case, it's 4a^2 MINUS 4ab + b^2

OpenStudy (inkyvoyd):

well \((x-y)^2=x^2-2xy+y^2\)

OpenStudy (anonymous):

so it would be (2a-b)^2!!!!! THANK YOU SO MUCH!!!

OpenStudy (inkyvoyd):

Sorry, I got it confused lol.

OpenStudy (inkyvoyd):

Do you get the second part?

OpenStudy (anonymous):

nope! lol

OpenStudy (anonymous):

the c squared?

OpenStudy (inkyvoyd):

Ok. Let's do it then ^^

OpenStudy (anonymous):

kaay(:

OpenStudy (inkyvoyd):

dyou follow? \(4a^2-4ab+b^2-c^2\) \(=(4a^2-4ab+b^2)-c^2\) \(=(2a-b)^2-c^2\) in LaTeX

OpenStudy (anonymous):

yes

OpenStudy (inkyvoyd):

Ok. remember (x+y)(x-y)=x^2-y^2 ?

OpenStudy (anonymous):

yes! so it would be (2a-b-c)(2a-b+c)!!!!!!!

OpenStudy (inkyvoyd):

Good :)

OpenStudy (anonymous):

so when there's a problem like this with a trinomial in it, how do i notice right off the bat?

OpenStudy (inkyvoyd):

Well, a tip is that \((ax+by)=(ax)^2+2*ax*by+(by)^2\) the trick is to realize that the coefficients are squared as well as the variables.

OpenStudy (anonymous):

alright! thanks so much(:

OpenStudy (inkyvoyd):

no problem :P

OpenStudy (anonymous):

:D